# Surg-3M: A Dataset and Foundation Model for Perception in Surgical Settings

### v1

@@ Abstract
Advancements in computer-assisted surgical procedures heavily rely on accurate visual data interpretation from camera systems used during surgeries. Traditional open-access datasets focusing on surgical procedures are often limited by their small size, typically consisting of fewer than 100 videos with less than 100K images. To address these constraints, a new dataset called Surg-3M has been compiled using a novel aggregation pipeline that collects high-resolution videos from online sources. Featuring an extensive collection of over 4K surgical videos and more than 3 million high-quality images from multiple procedure types, Surg-3M offers a comprehensive resource surpassing existing alternatives in size and scope, including two novel tasks. To demonstrate the effectiveness of this dataset, we present SurgFM, a self-supervised foundation model pretrained on Surg-3M that achieves impressive results in downstream tasks such as surgical phase recognition, action recognition, and tool presence detection. Combining key components from ConvNeXt, DINO, and an innovative augmented distillation method, SurgFM exhibits exceptional performance compared to specialist architectures across various benchmarks. Our experimental results show that SurgFM outperforms state-of-the-art models in multiple downstream tasks, including significant gains in surgical phase recognition (+8.9pp, +4.7pp, and +3.9pp of Jaccard in AutoLaparo, M2CAI16, and Cholec80), action recognition (+3.1pp of mAP in CholecT50) and tool presence detection (+4.6pp of mAP in Cholec80). Moreover, even when using only half of the data, SurgFM outperforms state-of-the-art models in AutoLaparo and achieves state-of-the-art performance in Cholec80. Both Surg-3M and SurgFM have significant potential to accelerate progress towards developing autonomous robotic surgery systems. Our dataset, models, and code are publicly available at https://github.com/visurg-ai/surg-3m.
@@ 1 Introduction
Contributions. We propose a novel data curation pipeline (Fig. 3) for collecting videos of selected surgical procedures, capturing exclusively surgical footage while discarding non-surgical content. Building upon this data curation pipeline, we present Surg-3M, a dataset for computer vision applications in surgery. Our Surg-3M dataset is large and diverse, containing 4K surgical videos and 3M images (Table 1). To the best of our knowledge, Surg-3M is the largest open-access surgical video dataset to date, 18×18$\times$18 × larger than the largest published surgical dataset [54]. To model the features of surgical video frames, we provide SurgFM, a pretrained model based on the ConvNeXt architecture [36] and DINO [8]. Additionally, we propose a novel augmented distillation method (Fig. 4) designed to enhance the local-to-global self-distillation mechanism described in  [8]. SurgFM outperformed the current state-of-the-art models (Fig. 1) on Cholec80 [54], M2CAI16 [54, 53], AutoLaparo [57], and CholecT50 [44]. Additionally, it helps to significantly reduce the annotation effort (Fig. 5). Building on our Surg-3M annotations, we introduce two new supervised learning applications: multi-label surgical procedure classification and surgery type classification. Leveraging our image foundation model, SurgFM, we propose SurgFM-Vid, a video classifier that achieves top performance on these new tasks, serving as a baseline for future research.
@@
Dataset Year # Procedure types # Videos # Frames
M2CAI16 [54] $2016$ 1 41 $95$K
Cholec80 [54] $2016$ 1 80 $183$K
CholecT50 [44] $2022$ 1 50 $101$K
AutoLaparo [57] $2022$ 1 21 $83$K
SAR-RARP50 [48] $2023$ 1 50 $163$K
GraSP [4] $2024$ $1$ $13$ $117$K
Surg-3M (ours) $\mathbf{2025}$ $\mathbf{35}$ $\mathbf{4194}$ $\mathbf{3}$M
@@ 3 Proposed Surg-3M Dataset
To develop a robust surgical foundation model, it is essential to create a comprehensive, large-scale dataset of diverse surgical images that covers the whole spectrum of minimally invasive surgical procedures executed within the operating room. To fulfill this need we propose Surg-3M, a dataset that contains 4K videos and 3M images corresponding to 35 different surgical procedures, including videos from robotic-assisted surgeries and non-robotic traditional laparoscopies (more details on these surgical techniques are provided in Sec. A of the supplementary material).
Ground truth and proposed new tasks. Each of the 4K surgical videos is annotated for two tasks: 1) multi-label procedure type classification, where the objective is to identify the procedure[s] performed in an intervention’s video from among 35 possible options (Fig. 2); and 2) binary surgery type classification, where the goal is to indicate whether a surgery is conducted robotically or non-robotically. The rationale for these annotations is two-fold. Firstly, it allows us to validate the comprehensiveness of our dataset by ensuring that it encompasses a broad spectrum of surgical procedures. Secondly, it facilitates the development of new methods that use this information as a key input (e.g., our SurgFM foundation model, video classification/retrieval, video similarity measurement, procedure-specific estimation of surgical remaining time).
Scenes in surgical videos unfold at a relatively slow pace, as surgeons must ensure patient safety. Consequently, our 3M image dataset was compiled by sampling our dataset of 4194 curated videos at 1 fps, which is standard practice in the field of surgical computer vision [54, 57, 44, 48, 53]. In the following sections we detail how we collected, curated, and annotated these videos. An overview of the curation pipeline is shown in Fig. 3.

### v2

@@ Abstract
Advancements in computer-assisted surgical procedures heavily rely on accurate visual data interpretation from camera systems used during surgeries. Traditional open-access datasets focusing on surgical procedures are often limited by their small size, typically consisting of fewer than 100 videos with less than 100K images. To address these constraints, a new dataset called Surg-3M has been compiled using a novel aggregation pipeline that collects high-resolution videos from online sources. Featuring an extensive collection of over 4K surgical videos totaling 938 hours of high-quality footage across multiple procedure types, Surg-3M offers a comprehensive resource surpassing existing alternatives in size and scope, including two novel tasks. To demonstrate the effectiveness of this dataset, we present SurgFM, a self-supervised foundation model pretrained on Surg-3M that achieves impressive results in downstream tasks such as surgical phase recognition, action recognition, and tool presence detection. Combining key components from ConvNeXt, DINO, and an innovative augmented distillation method, SurgFM exhibits exceptional performance compared to specialist architectures across various benchmarks. Our experimental results show that SurgFM outperforms state-of-the-art models in multiple downstream tasks, including significant gains in surgical phase recognition (+8.9pp, +4.7pp, and +3.9pp of Jaccard in AutoLaparo, M2CAI16, and Cholec80), action recognition (+3.1pp of mAP in CholecT50) and tool presence detection (+4.6pp of mAP in Cholec80). Moreover, even when using only half of the data, SurgFM outperforms state-of-the-art models in AutoLaparo and achieves state-of-the-art performance in Cholec80. Both Surg-3M and SurgFM have significant potential to accelerate progress towards developing autonomous robotic surgery systems. Our dataset, models, and code are publicly available at https://github.com/visurg-ai/surg-3m.
@@ 1 Introduction
Contributions. We propose a novel data curation pipeline (Fig. 3) for collecting videos of selected surgical procedures, capturing exclusively surgical footage while discarding non-surgical content. Building upon this data curation pipeline, we present Surg-3M, a dataset for computer vision applications in surgery. Our Surg-3M dataset is large and diverse, comprising more than 4K surgical videos totaling 938 hours of high-quality footage (Table 1). To the best of our knowledge, Surg-3M is the largest open-access surgical video dataset to date, 18×18$\times$18 × larger than the largest published surgical dataset [54]. To model the features of surgical video frames, we provide SurgFM, a pretrained model based on the ConvNeXt architecture [36] and DINO [8]. Additionally, we propose a novel augmented distillation method (Fig. 4) designed to enhance the local-to-global self-distillation mechanism described in  [8]. SurgFM outperformed the current state-of-the-art models (Fig. 1) on Cholec80 [54], M2CAI16 [54, 53], AutoLaparo [57], and CholecT50 [44]. Additionally, it helps to significantly reduce the annotation effort (Fig. 5). Building on our Surg-3M annotations, we introduce two new supervised learning applications: multi-label surgical procedure classification and surgery type classification. Leveraging our image foundation model, SurgFM, we propose SurgFM-Vid, a video classifier that achieves top performance on these new tasks, serving as a baseline for future research.
@@
Dataset # Procedures # Videos # Hours # Frames # Frames (1 fps)
M2CAI16 [54] 1 41 $26$ $2.4$M $95$K
Cholec80 [54] 1 80 $51$ $4.6$M $183$K
CholecT50 [44] 1 50 $28$ $2.5$M $101$K
AutoLaparo [57] 1 21 $23$ $2.1$M $83$K
SAR-RARP50 [48] 1 50 $45$ $9.8$M $163$K
GraSP [4] $1$ $13$ $33$ $3.5$M $117$K
Surg-3M (ours) $\mathbf{35}$ $\mathbf{4194}$ $\mathbf{938}$ $\mathbf{85}$M $\mathbf{3.4}$M
@@ 3 Proposed Surg-3M Dataset
To develop a robust surgical foundation model, it is essential to create a comprehensive, large-scale dataset of diverse surgical images that covers the whole spectrum of minimally invasive surgical procedures executed within the operating room. To fulfill this need we propose Surg-3M, a dataset comprising more than 4K videos totaling 938 hours of surgical footage covering 35 distinct procedures, including videos from robotic-assisted surgeries and non-robotic traditional laparoscopies (more details on these surgical techniques are provided in Sec. A of the supplementary material).
Ground truth and proposed new tasks. Each of the 4194 surgical videos is annotated for two tasks: 1) multi-label procedure type classification, where the objective is to identify the procedure[s] performed in an intervention’s video from among 35 possible options (Fig. 2); and 2) binary surgery type classification, where the goal is to indicate whether a surgery is conducted robotically or non-robotically. The rationale for these annotations is two-fold. Firstly, it allows us to validate the comprehensiveness of our dataset by ensuring that it encompasses a broad spectrum of surgical procedures. Secondly, it facilitates the development of new methods that use this information as a key input (e.g., our SurgFM foundation model, video classification/retrieval, video similarity measurement, procedure-specific estimation of surgical remaining time).
Scenes in surgical videos unfold at a relatively slow pace, as surgeons must ensure patient safety. Consequently, we extract frames from Surg-3M at 1 fps for our foundation model pretraining, following the standard practice in the field of surgical computer vision [54, 57, 44, 48, 53]. In the following sections we detail how we collected, curated, and annotated these videos. An overview of the curation pipeline is shown in Fig. 3.